### v2

@@ 3 Methodology
[…]
Figure 3: Different operationalizations of multimodality. The colored shapes are the detected communities, and the dashed boxes show on which network the community discovery algorithm is performed.
@@ 3.1 Preliminaries
The analytical framework for coordinated behavior detection exploiting network science-based methods can be structured in four steps: user selection, coordination network construction, network filtering, and community detection [1]. The user selection step identifies a subset $U^{\prime}\subseteq U$ of users based on criteria aligned with the analysis goals, often focusing on the most active accounts, since a small fraction of users generates the majority of activity on social platforms [54]. Common strategies include selecting super-producers, who post frequently [55, 29, 28], or super-spreaders, whose content is widely reshared [3, 56, 57]. Then, the second step involves the construction of the coordination network that in the most general case is a multiplex network $G(V,E,W,L)$ [58]. To build $G$, one must define the types of co-actions to consider (e.g., re-shares, mentions, follows, etc.). When multiple co-actions are used, $G$ is a multiplex network with $|L|$ layers, where $L$ is the set of layers, each corresponding to a selected co-action. However, the majority of existing works leverage a single co-action, i.e., the number of layers is $|L|=1$ and $G$ is a single layer network. Each layer $l\in L$ is a monomodal network in $G$ and is an undirected weighted graph $G^{l}(V^{l},E^{l},W^{l})$, where $V^{l}\subseteq U^{\prime}$, $E^{l}$ and $W^{l}$ respectively denote the set of nodes, edges, and weights of layer $l$. Given a layer $l$, an edge $e^{l}_{jk}$ is created if there exists a co-action of type $l$ between two users $(u_{j},u_{k})$. The edge weight $w^{l}_{jk}$ is obtained via a similarity function that computes pairwise user similarities. Network filtering is an optional but important step used to retain only meaningful coordination signals while reducing network size for more efficient analysis. A common strategy removes edges with weights below a threshold and the resulting isolated nodes [26, 30, 28, 51, 59, 44, 47]. Another approach filters based on the timing of co-actions, keeping only those occurring within specific time windows – either adjacent [60, 8, 19, 18] or overlapping [30, 52, 6, 28, 33]. The final step detects coordinated communities through a community discovery algorithm on the coordination network. Multiplex networks are either flattened prior to discovery [61, 48, 19] or analyzed using multiplex-specific algorithms [33, 58, 62].
@@ 3.2 Operationalizations of multimodality
Independent layers. This implementation detects coordinated behavior by identifying communities in each layer independently and then performing comparisons between the detected communities. So it leverages multiple monomodal approaches, focusing on each type of co-action independently. This is the most relaxed form of multimodality that completely loses the interplay between the different types of co-actions. For each modality $l\in L$, a community detection algorithm identifies a set of communities, similarly to the monomodal approach. In the rest of the work, we refer to this approach as INDI.
@@
Multiplex community detection. Multimodality is operationalized by performing a community discovery algorithm designed for multiplex networks. In the literature, several algorithms have been proposed for this purpose [58], with the advantage of leveraging inter-layer connections, which are implicit in multiplex networks. In such networks, two nodes representing the same user across different layers are implicitly linked, allowing the algorithm to capture multimodal coordination. Formally, a community detection algorithm is applied to the full multiplex network, yielding a set of communities that might span multiple layers. In the rest of the work, we refer to this approach as MULTI.
@@ 3.3.1 Overlapping communities
$o_{ij}^{AB}=\frac{2\cdot r_{ij}^{AB}\cdot r_{ji}^{AB}}{r_{ij}^{AB}+r_{ji}^{AB}}$ (1)
@@ 3.3.2 Lost, common, and gained communities
In this section, we analyze the overlap matrices computed in the previous step to classify communities as lost, common, or gained. To address the first research question, we examine whether the results of the INDI approaches differ or largely overlap, assessing the uniqueness of each modality. Similarly, we compare UNFL and MULTI against INDI. The classification of communities as lost, gained, or common is inherently perspective-dependent, as it relies on which approach is chosen as the reference. When comparing two community detection methods – such as a MONO and a multimodal one – the same set of unmatched communities may be considered either lost or gained depending on the direction of the comparison. In our analysis, we adopt the perspective of the multimodal approach as the reference, since our goal is to evaluate its contribution relative to the monomodal baselines. By doing so, we can properly interpret which communities are newly identified by the multimodal view, which ones are retained from the monomodal ones, and which are no longer detected. This choice of perspective ensures that the labeling captures the added value of multimodal integration. Hence, in the following analysis, we designate the multimodal approaches as the reference method, denoted by approach $B$. Instead, when comparing the INDI approaches, the choice of the reference method is less critical, although a MONO approach must still be selected as the baseline for consistency. For instance, in this work, we adopt the co-retweet modality as the reference.
Formally, the first step is the execution of the Hungarian algorithm [63], which provides the set of the couple of communities maximizing the sum of the harmonic mean overlap. Formally, let $\mathcal{P}=\{1,\dots,k\}\times\{1,\dots,k^{\prime}\}$ be the Cartesian product of the communities of the approach $A$ and $B$, the Hunagrian algorithm finds an assignment $\pi^{*}$ such that the total overlap is maximized, i.e., $\pi^{*}=\arg\max_{\pi\in\mathcal{P}}\sum_{(i,j)\in\pi}o_{ij}$, where $\pi^{*}\subseteq\mathcal{P}$ is a one-to-one matching. We must note that $\pi^{*}$ is a subset, since if $k>k^{\prime}$, not all communities of $C^{B}$ are matched. Each pair $(i,j)\in\pi^{*}$ represents a match between community $A_{i}$ and community $B_{j}$. Hence, we define the non-matching communities of approach $A$ as $A_{\text{unmatched}}=\left\{i\in\{1,\dots,k\}\;\middle|\;\nexists j\in\{1,\dots,k^{\prime}\}\text{ such that }(i,j)\in\pi^{*}\right\}$ and similarly, for approach $B$, $B_{\text{unmatched}}=\left\{j\in\{1,\dots,k^{\prime}\}\;\middle|\;\nexists i\in\{1,\dots,k\}\text{ such that }(i,j)\in\pi^{*}\right\}$.
@@
$\lambda^{lost}(C_{i}^{A},O)=\begin{cases}\text{common}&\text{if }\exists j\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}\geq\theta\\ \text{lost}&\text{if }\exists j\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}<\theta\\ \text{lost}&\text{if }i\in A_{\text{unmatched}}\end{cases}$ (3)
$\lambda^{gain}(C_{j}^{B},O)=\begin{cases}\text{common}&\text{if }\exists i\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}\geq\theta\\ \text{gained}&\text{if }\exists i\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}<\theta\\ \text{gained}&\text{if }j\in B_{\text{unmatched}}\end{cases}$ (4)
@@ 3.3.3 Lost, common, and gained nodes
$\lambda^{node}(V,C_{i}^{A},C_{j}^{B},O)=\begin{cases}\text{common}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{i}^{A}\cap C_{j}^{B}\\ \text{lost}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{i}^{A}\setminus C_{j}^{B}\\ \text{gained}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{j}^{B}\setminus C_{i}^{A}\\ \text{lost}&\text{if }\exists i\in A_{\text{unmatched}}\text{ such that }v\in C_{i}^{A}\\ \text{gained}&\text{if }\exists j\in B_{\text{unmatched}}\text{ such that }v\in C_{j}^{B}\end{cases}$ (5)
@@ 4 Experimental setting
In this section, we present the dataset utilized for the experiments, detail the preprocessing steps applied, and describe the implementation choices adopted for the method. The proposed method can be applied to any dataset that contains multiple types of social media actions, such as sharing, replying, or publishing posts with hashtags. In this study, we use the dataset from Nizzoli et al. [3], collected via Twitter Streaming APIs during the 2019 UK General Election. The collection spans from 12 November to 12 December 2019, covering the month leading up to the election. It includes all tweets mentioning at least one of the most popular election-related hashtags, as well as all tweets from the official accounts of the two main parties and their leaders, along with the interactions (retweets and replies) those tweets received.
@@ 4.1 Preprocessing
Some initial decisions are needed to be taken to define the input dataset for the proposed method. We filter out the hashtags and mentions adopted for the data collection, as they would alter the results of the analysis, since we are searching for common actions. So the following hashtags have been filtered: GE2019, GeneralElection19, GeneralElection19, GeneralElection2019, VoteLabour, VoteLabour2019, ForTheMany, ForTheManyNotTheFew, ChangeIsComing, RealChange, VoteConservative, VoteConservative2019, BackBoris, GetBrexitDone. The following mentions have been filtered: jeremycorbyn, UKLabour, BorisJohnson, Conservatives.
In the literature, the domain is often considered instead of the full URL, as it provides a more generalized and meaningful representation of the source [64, 65, 59]. Following this approach, we extracted the domain from the URLs to focus on this higher-level identifier. Subsequently, the most frequently shared URLs are manually reviewed to filter out [^1] The complete list of domain URLs removed is: a) twitter.com, b) cards.twitter.com, c) youtube.com, d) instagram.com, e) open.spotify.com, f) google.com, g) reddit.com, h) play.google.com, i) bing.com, j) google.co.uk. those that, based on qualitative assessment within the specific context of the dataset, are not considered plausible targets of coordination, aligning with established methodologies and enhancing the interpretability of the analysis [60, 65, 26].
@@ 4.2 Implementation choices
Network-science method. In the user selection step, we aim to maximize the number of users while ensuring sufficient overlap across different actions, without incurring high computational costs. To this end, for each action type $l$, we select the top 5% most active users, denoted as $U^{\prime}_{l}$, and define the final user set $U^{\prime}$ as the union of these subsets. While this does not guarantee that each user is active in all modalities, it enables us to retain more users compared to prior studies [3, 56, 57], while ensuring substantial inter-layer overlap and a shared user base for multimodal comparisons.
@@
The second step is the coordination network construction. We focus on the most common actions (co-actions) on Twitter/X, i.e., co-retweet (RTW), co-reply (RPL), co-mention (MEN), co-hashtag (HST), co-URL (URL). We build a multiplex network with five layers, each corresponding to a different co-action. Every user is represented by a vector capturing their activity: retweeted post IDs, replied-to tweet IDs, mentioned user IDs, hashtags, and shared URLs. In each layer, we calculate the edge weight between pairs of users using the cosine similarity of their TF-IDF-weighted vectors [1]. This approach reduces the influence of widely shared or viral content and emphasizes the significance of less common items [3, 57]. We also compute, the absolute overlapping between the vectors representing the users to have a measure of the number of common actions generating the edge. This measure is useful to understand the significance of the edge.
@@
Moreover, in this work, we consider an overlapping time window, evenly distributed in time, of 6 hours with a shift of 5 hour and consequently a 1-hour overlap. This choice is is motivated by the fact that we want to capture both spontaneous and malicious coordination, without focusing on a specific type of behavior. In this way, according to the period of reference (one month), we obtain around 147 time windows (the precise number can vary according to the action). In this work, we build a coordination network for each time window and then we merge them into a single network by averaging the weights of edges across time windows and summing those that appeared in multiple windows. The final coordination network is a multiplex network with 5 layers, one for each type of action.
Network filtering step aims to filter out insignificant edges to highlight meaningful coordination, while preserving as much user overlap across layers as possible. First, edges generated by few common actions are removed, using action-specific thresholds chosen to retain up to $20,000$ nodes per layer, balancing coverage and computational cost. Then, a second filter is applied to edge weights, using the median weight in each layer as a threshold to retain a representative set of edges. This two-step process ensures both relevance and scalability, as detailed in Table 1.
@@
Table 1: Number of nodes and edges in the coordination network before and after the filtering steps. The thresholds are set on the number of common actions generating the edge ($th_{a}$) and on the edge weights ($th_{w}$).
@@
Multimodal operationalizations. The implementations choices for the multimodal operationalizations described in Section 3.2 are the following:
@@
  • INDI is implemented by performing Louvain algorithm [66] on each layer independently. The resolution parameter is set to 1, allowing the detection of communities of varying sizes.
@@
  • UNFL is implemented by flattening the multiplex network into a single-layer network, with 3 different strategies of weight aggregation: 1. UNFL(nw): not weighted edges [31, 26, 30], 2. UNFL(ec): weight equal to the number of layers in which the edge is present, 3. UNFL(sum): weight equal to the sum of the weights of the edges in the different layers [2]. We then apply the Louvain algorithm on the flattened network, setting the resolution parameter to 1.
@@
  • MULTI is implemented by applying the Generalized Louvain algorithm [67] on the multiplex network. This method adapts the Louvain algorithm for multiplex networks and it maximizes modularity across all layers, accounting for both intra-layer and inter-layer structures. A key feature is the coupling parameter $\omega$, which regulates the weight of inter-layer connections. This approach enables communities to span layers without requiring presence in all of them, offering a balanced form of multimodality. In this study, $\omega$ is fixed to enable fair comparison across different multimodal operationalizations. We set the resolution parameter $\gamma=1$ and the coupling parameter $\omega=0.1$.
@@
  • INTFL is implemented by flattening the multiplex network into a single-layer network, considering the intersection of the edges and nodes of the different layers and the weight equal to the sum of the weights of the edges in the different layers. We apply the Louvain algorithm on the flattened network, setting the resolution parameter to 1.

### v3

@@ 3 Methodology
Figure 3: Different operationalizations of multimodality. The colored shapes are the detected communities, and the dashed boxes show on which network the community detection algorithm is performed.
@@ 3.1 Preliminaries
The analytical framework for coordinated behavior detection exploiting network science-based methods can be structured in four steps: user selection, coordination network construction, network filtering, and community detection [1]. The user selection step identifies a subset $U^{\prime}\subseteq U$ of users based on criteria aligned with the analysis goals, often focusing on the most active accounts, since a small fraction of users generates the majority of activity on social platforms [36]. Common strategies include selecting super-producers, who post frequently [18, 17], or super-spreaders, whose content is widely reshared [3, 37]. Then, the second step involves the construction of the coordination network that in the most general case is a multiplex network $G(V,E,W,L)$ [38]. To build $G$, one must define the types of co-actions to consider (e.g., re-shares, mentions, follows, etc.). When multiple types of co-actions are used, $G$ is a multiplex network with $|L|$ layers, where $L$ is the set of layers, each corresponding to a selected co-action. However, the majority of existing works leverage a single co-action, i.e., the number of layers is $|L|=1$ and $G$ is a single-layer network. Each layer $l\in L$ is a monomodal network in $G$ and is an undirected weighted graph $G^{l}(V^{l},E^{l},W^{l})$, where $V^{l}\subseteq U^{\prime}$, $E^{l}$ and $W^{l}$ respectively denote the set of nodes, edges, and weights of layer $l$. Given a layer $l$, an edge $e^{l}_{jk}$ is created if there exists a co-action of type $l$ between two users $(u_{j},u_{k})$. The edge weight $w^{l}_{jk}$ is obtained via a similarity function that computes pairwise user similarities. Network filtering is an optional but important step used to retain only meaningful coordination signals while reducing network size for more efficient analysis. A common strategy removes edges with weights below a threshold and the resulting isolated nodes [15, 33]. Another approach filters based on the timing of co-actions, keeping only those occurring within specific time windows – either adjacent [7, 11] or overlapping [5, 17]. The final step detects coordinated communities through a community detection algorithm on the coordination network. Multiplex networks are either flattened prior to detection [39, 11] or analyzed using multiplex-specific algorithms [38].
@@ 3.2 Operationalizations of multimodality
Independent layers. This implementation detects coordinated behavior by identifying communities in each layer independently and then performing comparisons between the detected communities. So it leverages multiple monomodal approaches, focusing on each type of co-action independently. This is the most relaxed form of multimodality that completely loses the interplay between the different types of co-actions. For each modality $l\in L$, a community detection algorithm identifies a set of communities, similarly to the monomodal approach. In the rest of the work, we refer to this approach as INDE.
@@
Multiplex community detection. Multimodality is operationalized by performing a community detection algorithm designed for multiplex networks. In the literature, several algorithms have been proposed for this purpose [38, 40], with the advantage of leveraging inter-layer connections, which are implicit in multiplex networks. In such networks, two nodes representing the same user across different layers are implicitly linked, allowing the algorithm to capture multimodal coordination. Formally, a community detection algorithm is applied to the full multiplex network, yielding a set of communities that might span multiple layers. In the rest of the work, we refer to this approach as MULTI.
@@ 3.3.1 Overlapping communities
$o_{ij}^{AB}=\frac{2\cdot r_{ij}^{AB}\cdot r_{ji}^{AB}}{r_{ij}^{AB}+r_{ji}^{AB}}~~~~~~,$ (1)
@@ 3.3.2 Lost, common, and gained communities
In this section, we analyze the overlap matrices computed in the previous step to classify communities as lost, common, or gained. To address the first research question, we examine whether the results of the INDE approaches differ or largely overlap, assessing the uniqueness of each modality. Similarly, we compare UNFL and MULTI against INDE. The classification of communities as lost, gained, or common is inherently perspective-dependent, as it relies on which approach is chosen as the reference. When comparing two community detection methods – such as a MONO and a multimodal one – the same set of unmatched communities may be considered either lost or gained depending on the direction of the comparison. In our analysis, we adopt the perspective of the multimodal approach as the reference, since our goal is to evaluate its contribution relative to the monomodal baselines. By doing so, we can properly interpret which communities are newly identified by the multimodal view, which ones are retained from the monomodal ones, and which are no longer detected. This choice of perspective ensures that the labeling captures the added value of multimodal integration. Hence, in the following analysis, we designate the multimodal approaches as the reference method, denoted by approach $B$. Instead, when comparing the INDE approaches, the choice of the reference method is less critical, although a MONO approach must still be selected as the baseline for consistency. For instance, in this work, we adopt the co-retweet modality as the reference.
Formally, the first step is the execution of the Hungarian algorithm [41], which provides the subset of the couple of communities maximizing the sum of the harmonic mean overlap. Formally, let $\mathcal{P}=\{1,\dots,k\}\times\{1,\dots,k^{\prime}\}$ be the Cartesian product of the communities of the approach $A$ and $B$, the Hunagrian algorithm finds an assignment $\pi^{*}$ such that the total overlap is maximized, i.e., $\pi^{*}=\arg\max_{\pi\in\mathcal{P}}\sum_{(i,j)\in\pi}o_{ij}$, where $\pi^{*}\subseteq\mathcal{P}$ is a one-to-one matching. We must note that $\pi^{*}$ is a subset, since if $k>k^{\prime}$, not all communities of $C^{B}$ are matched. Each pair $(i,j)\in\pi^{*}$ represents a match between community $A_{i}$ and community $B_{j}$. Hence, we define the non-matching communities of approach $A$ as $A_{\text{unmatched}}=\left\{i\in\{1,\dots,k\}\;\middle|\;\nexists j\in\{1,\dots,k^{\prime}\}\text{ such that }(i,j)\in\pi^{*}\right\}$ and similarly, for approach $B$, $B_{\text{unmatched}}=\left\{j\in\{1,\dots,k^{\prime}\}\;\middle|\;\nexists i\in\{1,\dots,k\}\text{ such that }(i,j)\in\pi^{*}\right\}$.
@@
[…]
@@ 3.3.3 Lost, common, and gained nodes
[…]
@@ 4 Experimental setting
In this section, we present the dataset utilized for the experiments, detail the preprocessing steps applied, and describe the implementation choices adopted for the method. [^1] https://zenodo.org/records/17709815
4.1 Data
The proposed method can be applied to any dataset that contains multiple types of social media actions, such as sharing, replying, or publishing posts with hashtags. In this study, we use two datasets: UK from Nizzoli et al. [3] and IORussia from Seckin et al. [42].
UK dataset. UK dataset is collected via Twitter Streaming API during the 2019 UK General Election. The collection spans from 12 November to 12 December 2019, covering the month leading up to the election. It includes all tweets mentioning at least one of the most popular election-related hashtags, as well as all tweets from the official accounts of the two main parties and their leaders, along with the interactions (retweets and replies) those tweets received. The final dataset includes a total of $11,264,820$ tweets published by $1,179,659$ distinct users.
This dataset is publicly available for research purposes [^2] https://doi.org/10.5281/zenodo.4647893 and has been widely adopted in the study of coordinated online behavior [20, 3, 14]. It has been shown to include several coordinated communities, making it an appropriate benchmark for our analysis. The data contains user interactions, retweet relationships, and temporal information. Previous research has primarily analyzed this dataset through the co-retweet network. In contrast, we extend the analysis to a multimodal representation that integrates multiple types of user interactions, enabling a deeper detection of coordination patterns beyond retweet activity alone.
IORussia dataset. The IORussia dataset is part of a collection of 26 Twitter campaigns involving information operations (IOs), each containing posts verified by a social media platform as part of coordinated activity, along with accounts that discussed similar topics within the same time frames (control data). In this work, we focus on the largest campaign in terms of users, namely “Russia_1”. Posts and user information belonging to the IO campaign were collected from a public archive of state-sponsored information operations, made available through the platform’s transparency website. The authors then retrieved all hashtags used by IO accounts within each campaign to identify control accounts (ctrl) discussing similar topics. These hashtags were used as search queries via the platform’s API to collect accounts that posted on the same dates and used the same hashtags as the IO accounts. Finally, the daily timelines of the control accounts were reconstructed by extracting up to 100 messages posted on the same dates as the IO accounts. The IORussia dataset spans the period from 2009 to 2018 and includes, respectively, $3,293$ IO accounts and $31,303$ control accounts, with $1,826,344$ and $1,823,859$ posts. This dataset is publicly available for research purposes. [^3] https://zenodo.org/records/14189193
The IORussia dataset presents several features that make it particularly interesting for our analysis. First, it provides two labeled groups that can be leveraged as ground-truth for coordinated behavior detection: IO accounts are considered coordinated, while control accounts are labeled as non-coordinated. Although this labeling is an approximation – since IO campaigns typically rely on a coordinated core of malicious accounts to be effective [43] – it is valuable in a domain where reliable ground-truth is scarce. This labeling strategy has already been adopted in previous studies [7, 10]. Secondly, compared to the UK dataset, which covers approximately one month, the IORussia dataset spans nine years. Conducting experiments on datasets with such different temporal scales allows for an analysis of coordination patterns over both short and long time frames. While most studies in the literature focus on short- or medium-term coordination, the examination of long-term coordination is equally meaningful, as it captures the dynamic and evolving nature of coordinated behavior over time.
@@ 4.2 Preprocessing
Some data preprocessing steps are necessary for preparing the datasets for the proposed methods. In UK dataset, we filter out the hashtags and mentions adopted for the data collection, as they would alter the results of the analysis, since we are searching for common actions. So the following hashtags have been filtered out: GE2019, GeneralElection19, GeneralElection19, GeneralElection2019, VoteLabour, VoteLabour2019, ForTheMany, ForTheManyNotTheFew, ChangeIsComing, RealChange, VoteConservative, VoteConservative2019, BackBoris, GetBrexitDone. Likewise, the following mentions have been filtered out: jeremycorbyn, UKLabour, BorisJohnson, Conservatives. Instead, for the IORussia dataset, information about the specific hashtags used by IO accounts to identify and collect control accounts is not publicly available.
In the literature, the domain is often considered instead of the full URL, as it provides a more generalized and meaningful representation of the source [44]. Following this approach, for UK dataset we extracted the domain from the URLs to focus on this higher-level identifier. Subsequently, the most frequently shared URLs are manually reviewed to filter out [^4] The complete list of domain URLs removed is: a) twitter.com, b) cards.twitter.com, c) youtube.com, d) instagram.com, e) open.spotify.com, f) google.com, g) reddit.com, h) play.google.com, i) bing.com, j) google.co.uk. those that, based on qualitative assessment within the specific context of the dataset, are not considered plausible targets of coordination, aligning with established methodologies and enhancing the interpretability of the analysis [15]. In the IORussia dataset, URLs have been hashed for privacy reasons, preventing the extraction of their domains. Therefore, in this case study, we rely on the complete hashed URL strings instead.
@@ 4.3 Implementation choices
Network-science method. In the user selection step, we aim to maximize the number of users while ensuring sufficient overlap across different actions, without incurring high computational costs. To this end, for UK dataset, for each action type $l$, we select the top 5% most active users, denoted as $U^{\prime}_{l}$, and define the final user set $U^{\prime}$ as the union of these subsets. While this does not guarantee that each user is active in all modalities, it enables us to retain more users compared to prior studies [3, 37], while ensuring substantial inter-layer overlap and a shared user base for multimodal comparisons. Conversely, the IORussia dataset is smaller in scale and computationally feasible to process. Thus, we retained all users for the analysis of such dataset.
@@
The second step is the coordination network construction. We focus on the most common actions (co-actions) on Twitter/X – such as, co-retweet (RTW), co-reply (RPL), co-mention (MEN), co-hashtag (HST), co-URL (URL). For both datasets, we build a multiplex network with five layers, each corresponding to a different co-action. Every user is represented by a vector capturing their activity: retweeted post IDs, replied-to tweet IDs, mentioned user IDs, hashtags, and shared URLs. In each layer, we calculate the edge weight between pairs of users using the cosine similarity of their TF-IDF-weighted vectors [1]. This approach reduces the influence of widely shared or viral content and emphasizes the significance of less common items [3, 37]. We also compute the absolute overlapping between the vectors representing the users to have a measure of the number of common actions generating the edge. This measure is useful to understand the significance of the edge.
@@
Moreover, in this work, for both datasets, we consider an overlapping time window, evenly distributed in time. Specifically, for UK of 6 hours with a shift of 5 hours and consequently a 1-hour overlap. This choice is motivated by the fact that we want to capture both spontaneous and malicious coordination, without focusing on a specific type of behavior. In this way, according to the period of reference (one month), we obtain around 147 time windows (the precise number can vary according to the action). In the case of the IORussia dataset, we applied a 7-day time window with a 6-day shift, leading to a 1-day overlap and approximately 470 time windows. Given that this dataset covers 9 years, we used broader temporal windows to capture longer-term coordination dynamics. In this work, we build a coordination network for each time window and then we merge them into a single network by averaging the weights of edges across time windows and summing those that appeared in multiple windows. The final coordination network is a multiplex network with 5 layers, one for each type of action.
@@
Table 1: Number of nodes and edges in the UK coordination network before and after the filtering steps. The thresholds are set on the number of common actions generating the edge ($th_{a}$) and on the edge weights ($th_{w}$).
@@
Table 2: Number of nodes and edges in the IORussia coordination network before and after the filtering steps. The thresholds are set on the number of common actions generating the edge ($th_{a}$) and on the edge weights ($th_{w}$). The last three columns report the percentages of nodes filtered for the overall dataset and separately for the control (ctrl) and coordinated (coord) groups.
no filter filtered on actions filtered on weight filteringRates (%)
layer nNodes nEdges $\boldsymbol{th_{a}}$ nNodes nEdges $\boldsymbol{th_{w}}$ nNodes nEdges %overall %ctrl %coord
RTW $8,090$ $104,478$ 2 $4,604$ $55,288$ 0.14 $3,397$ $27,566$ 76.85 81.63 41.21
RPL $1,409$ $5,520$ 2 $436$ $1,652$ 0.58 $203$ $826$ 97.66 99.57 66.40
URL $19,603$ $3,515,318$ 2 $2,533$ $11,393$ 0.27 $2,327$ $5,698$ 88.14 89.51 77.35
MEN $11,263$ $179,360$ 2 $6,920$ $98,862$ 0.12 $4,874$ $49,478$ 74.02 78.68 34.71
HST $21,038$ $1,358,160$ 7 $4,367$ $67,102$ 0.25 $3,169$ $33,512$ 85.77 90.32 43.14
Network filtering step aims to filter out insignificant edges to highlight meaningful coordination, while preserving as much user overlap across layers as possible. For UK dataset, first, edges generated by few common actions are removed, using action-specific thresholds chosen to retain up to $20,000$ nodes per layer, balancing coverage and computational cost. Then, a second filter is applied to edge weights, using the median weight in each layer as a threshold to retain a representative set of edges. This two-step process ensures both relevance and scalability, as detailed in Table 1.
Similarly, for the IORussia dataset, we applied the same two filtering procedures. In the first step, which relies on common actions, a lower threshold was adopted to account for the smaller network scale and to preserve an adequate number of nodes. For the second step, based on edge weights, we used the same criterion of UK, applying the median weight within each layer as the filtering threshold. Table 2 summarizes the details of the IORussia network filtering. Compared to the UK dataset, it additionally reports the percentages of nodes filtered for the overall dataset and separately for the control (ctrl) and coordinated (coord) groups. We observe that the filtering process removes a larger proportion of control users than coordinated users.
@@
Multimodal operationalizations. For all the multimodal operationalizations, we adopted two base methods, namely Louvain [45] and Infomap [46]. The implementation choices for the multimodal operationalizations described in Section 3.2 are the following:
@@
  • INDE is implemented by performing Louvain [45] and Infomap [46] algorithms on each layer independently. For Louvain, the resolution parameter is set to 1.
@@
  • UNFL is implemented by flattening the multiplex network into a single-layer network, with 3 different strategies of weight aggregation:
    • UNFL (nw): unweighted edges [19, 15, 31],
    • UNFL (ec): edge weight equal to the number of layers in which the edge is present,
    • UNFL (sum): edge weight equal to the sum of the weights of the edges in the different layers [2].
    We then apply the Louvain (setting the resolution parameter to 1) and Infomap algorithms on the flattened network.
@@
  • MULTI is implemented by applying the Generalized Louvain [47] and Generalized Infomap [40] algorithms on the multiplex network. The first one adapts Louvain for multiplex networks and maximizes modularity across all layers, accounting for both intra-layer and inter-layer structures. A key feature is the coupling parameter $\omega$, which regulates the weight of inter-layer connections. This approach enables communities to span layers without requiring presence in all of them, offering a balanced form of multimodality. In this study, $\omega$ is fixed to enable fair comparison across different multimodal operationalizations. We set the resolution parameter $\gamma=1$ and the coupling parameter $\omega=0.1$. Similarly, Generalized Infomap adapts Infomap for multiplex networks. In this case, the weight of inter-layer connections is set to the default value $w_{inter}=0.15$.
@@
  • INTFL is implemented by flattening the multiplex network into a single-layer network, considering the intersection of the edges and nodes of the different layers and the weight equal to the sum of the weights of the edges in the different layers. We apply Louvain (setting the resolution parameter to 1) and Infomap on the flattened network.